OpenStudy (anonymous):

Solve the equation. l6x+4l=34

OpenStudy (saifoo.khan):

6x+4 = 34 or 6x+4 = -34

OpenStudy (saifoo.khan):

6x+4 = 34 or 6x+4 = -34 6x = 30 or 6x = -34-8 x = 5 or x = 19/3

OpenStudy (anonymous):

Do I need both answers? Like for a test? Or only one is necessary?

OpenStudy (saifoo.khan):

Oh Yes, both!

OpenStudy (anonymous):

Thank you! Is it the same for inequalities with absolute value in it?

OpenStudy (saifoo.khan):

No, in equality, the signs interchanges.

OpenStudy (anonymous):

Wait, so the negative stays a negative if it's in an inequality? Like for this one...... l2x-3l(is greater than or equal to) 1